OpenStudy (happykiddo):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (cwrw238):

this is basically pythagoras theorem AB = sqrt(4-(-4)^2 + (3--3)^2)

OpenStudy (cwrw238):

sqrt (difference of x values squared + difference in y values squared)

OpenStudy (cwrw238):

here it is sqrt(8^2 + 6^2) = sqrt (64+36) = sqrt(100)

OpenStudy (happykiddo):

its 10 right

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

10

OpenStudy (happykiddo):

thank you :D